# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ VI AE based Communication systems
The main goal of a communication system, acting as an AE, is to find a robust latent space representation $w$ of the input data in the transmitter to counteract any impairments introduced by the channel. In particular, through an end-to-end training, the AE learns to effectively reconstruct the input data at the receiver, considering the presence of channel impairments. During the training process, the encoder component of the AE acquires resilient symbol sequence representations for all incoming data. Once training is complete, both the transmitter and the receiver can function independently, using their respective layer dimensions and weights. The remainder of this section delves into the various aspects of employing AEs in communication systems, encompassing the two AE approaches, activation functions, loss functions, physical constraints, and AE architectural variants.
VI-A Symbol-Wise and Bit-Wise AE
Symbol-wise AEs and Bit-Wise AEs are two different approaches to designing a communication system. Their primary difference lies in how they handle the mapping of symbols or bits from input messages to effectively determine the learnt constellation. This may directly impact system performance characteristics such as bit error rate (BER), symbol error rate (SER), and computational complexity. The learnt constellations of a symbol-wise AE model do not account for a bit-labelling scheme. Therefore, determining the optimal labelling scheme for the learnt constellations is a key challenge. To address this, a bit-wise AE approach offers a more direct optimisation path. To optimise geometric constellation shaping (GCS), symbol-wise AE can be trained for that, while bit-wise AE can be trained jointly for GCS and bit labelling [21]. In both symbol-wise AE and bit-wise AE, the input to the encoder AE can be pre-processed to one hot encoded.
@@ VI-A Symbol-Wise and Bit-Wise AE
Fig. 4 shows three different end-to-end learning architectures based on AE. Regression problems can be used in AEs for end-to-end learning of communication systems. By setting the output layer of the AE to a linear regression layer as shown in Fig. 4 (a), the AE can be trained to predict continuous values for new inputs[22, 23]. This results in complex output data consisting of both in-phase (I) and quadrature (Q) components. This data is then fed into a demapper for symbol detection and subsequently a hard decision circuit to recover the original bit sequence. The AE encoder processes complex baseband symbols ($x\in\mathbb{C}$) from modulation formats such as quadrature phase-shift keying (QPSK), or M-quadrature amplitude modulation ($M$-QAM). The encoder, parametrised by $\theta_{t}$, generates resilient latent representations ($w\in\mathbb{C}$) by mapping $f_{\theta_{t}}:x\mapsto w$. These representations are then passed through a channel model, where they undergo impairments. The decoder, parameterized by $\theta_{r}$, aims to reconstruct the original symbol from the channel-corrupted output ($y\in\mathbb{C}$) by mapping $f_{\theta_{r}}:y\mapsto\hat{x}$, where $\hat{x}\in\mathbb{C}$. As it tries to predict the data of a numerical symbol, the Mean Squared Error (MSE) is a suitable loss function; see Section VI-C. Since NNs cannot directly process complex data, the $N$ complex input data must be converted to $2N$ real-valued representation for both the encoder and decoder by implementing a $\mathbb{C}2\mathbb{R}$ layer. The resulting real-valued output is then converted back to complex data by using a $\mathbb{R}2\mathbb{C}$ layer.
@@
Alternatively, as shown in Fig. 4 (b), a symbol-wise approach based on minimising categorical cross-entropy (CCE) can be fed with a message $a\in\{0,1,...,M-1\}$, $M$ denotes the number of different symbols as $M=2^{m}$ where $m$ is the number of bits per symbol. Thus, the encoder then processes the mapping to complex data as $f_{\theta_{t}}:\{0,1,...,M-1\}\mapsto w$, i.e. $f_{\theta_{t}}:a\mapsto w$. Another key distinction lies in the output layer, the Softmax activation layer, which is used to obtain a probability vector of length $M$ which corresponds to the probabilities of possible $M$ messages, $p_{\theta_{r}}(a|y)$ [21]. This represents the estimated posterior distribution of the transmitted signal given the received signal. The constellation obtained through symbol-wise training is not optimised for bit-labelled. Bit-labelled indicates that each constellation point must be uniquely mapped to a specific bit vector. To enable joint constellation optimisation and bit labelling, a bit-wise AE can be used, as shown in Fig. 4 (c), can be used. $m$ length input bit vector $m$ length $m$ length $m$ length $b$ can be fed to the encoder to perform the mapping $f_{\theta_{t}}:b\mapsto w$. The AE output is used directly without applying an activation function. These output values are called logits ($l$) which is a raw data and represents un-normalised probabilities of the data belonging to a specific class. Logits are normally suitable for classification tasks. Therefore, the loss function of BCE is suitable for such an approach. The Sigmoid function outputs a probability for each bit independently, which is applied to the decoder logits $l$ This allows for the direct computation of the estimated posterior probabilities $p_{\theta_{r}}(b|y)$ for each bit, facilitating the use of log-likelihood ratios (LLRs) in the analysis [21]. The minimisation binary cross-entropy (BCC) function can then be used to optimise the AE, more details are given in Section VI-C.
VI-B Activation Functions
@@
In the context of using AE for end-to-end learning in communication systems, a list of possible activation functions can be seen in Table II. The ReLU is widely used in hidden layers of encoder and decoder parts of the AE model. The ELU activation function is a non-saturating function that accelerates training while mitigating the risk of gradient explosion or vanishing gradient problems. The existing literature commonly employs ReLU and/or ELU activation functions in the hidden layer of AE-based communication systems. For example, [24] initially employed ReLU in their AE model but later switched to ELU in [25] for improved results. Therefore, it is highly recommended to investigate both in order to determine the optimal choice. Additionally, further research is needed to determine the combination use of both in the AE model, for example, one for the encoder and the other for the decoder. However, the LeakyReLU activation function allows for a small, non-zero gradient when the input is negative, which helps mitigate the ”dying ReLU” problem often encountered with standard ReLU activations. This ensures a small gradient when the unit is inactive, which helps to keep the learning process alive [26]. Tanh activation has zero-centred output, which can make optimisation easier. It is also useful in hidden layers to model complex relationships. In [27], Tanh activation is used to extract the feature. In [28], it is used in all the 1D convolutional layers. The scaled exponential linear unit (SELU) can induce self-normalisation, which means that the activation automatically converges to zero mean and unit variance, as can be seen in Table II where $\lambda\approx 1.0507$ and $\alpha\approx 1.6733$. This helps maintain the stability of the model during training, which is crucial for AEs that require consistent encoding and decoding [29]. It is suited for the hidden layer of AEs to leverage its self-normalisation, which can lead to faster convergence. More research is required to investigate the potential benefits of alternative activation functions, such as Swish, for better gradient flow in the hidden layers. Finally, Softmax and Sigmoid activation layers are typically used in the output layer for classification tasks. The first can be used for symbol detection, while the second can be used for binary detection.
TABLE II: Common Activation Functions in AE-based communications
Name Function Output Range
ReLU $\text{ReLU}(x)=\max(0,x)$ [0, $\infty$)
ELU $\text{ELU}(x)=\begin{cases}x&\text{if }x\geq 0\\ \alpha(e^{x}-1)&\text{if }x<0\end{cases}$ ($-\alpha$, $\infty$)
Leaky ReLU Leaky $\text{Leaky ReLU}(x)=\begin{cases}x&\text{if }x\geq 0\\ \alpha x&\text{if }x<0\end{cases}$ ($\infty$- ∞, $\infty$)
Tanh $\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ (--1, 1)
SELU $SELU(x)=\begin{cases}\lambda x&\text{if }x>0\\ \lambda\alpha(e^{x}-1)&\text{if }x\leq 0\end{cases}$ [--1.758, $\infty$)
Swish $\text{Swish}(x)=x\cdot\sigma(x)$ ($\infty$- ∞, $\infty$)
Softmax $\text{Softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$ (0, 1)
Sigmoid $\sigma(x)=\frac{1}{1+e^{-x}}$ (0, 1)
VI-C Loss Functions
In the context of communication systems, there are two variant ways of dealing with the communication problem, either as a regression problem or a categorical problem [30]. A regression problem in the context of machine learning and communication systems typically involves predicting continuous or numerical data based on the input data. By setting the output layer of the AE to a regression layer (such as a linear layer), the AE can be trained to reconstruct a continuous target value, which can be used effectively in communication systems.
In contrast, in classification problems, the predicting data is a categorical or discrete class. End-to-end AE learning for communication systems using a classification problem involves training the AE to recognise and classify different constellation diagrams such as QPSK, 16-QAM, and 64-QAM, based on the structure and shape of the diagrams. This can be done by setting the output layer of the AE to a classification layer (such as the sigmod activation layer) and compute the cross-entropy. This way can be useful in applications such as the evaluation of symbol error rate and the evaluation of communication system performance.
VI-C1 Mean Squared Error (MSE)
@@
It serves as a loss function during the training phase, which is determined by computing the mean squared deviations between the predicted values and the actual ground truth values [31]. The general expression of the MSE is
@@
Gradient-based learning [32, 33] is employed to find optimal parameters $\theta=\{\theta_{t},\theta_{r}\}$ that minimize the loss function $\mathcal{L}_{MSE}(\theta)$ as
@@
where $\beta$ represents the learning rate, and the Stochastic Gradient Descent, such as Adam optimiser, enhances convergence.
VI-C2 Cross-Entropy (CE)
CE is a widely used metric for training classification models since it is differentiable, which is crucial for gradient-based optimisation algorithms[34, 35]. It is a quantitative metric in information theory that is used to measure the divergence between two probability distributions. This means that it measures the dissimilarity between two probability distributions. In the context of neural networks, one distribution is the output of the predicted probability distribution by the network, and the other is the true (target) distribution. Categorical cross-entropy (CCE) and sparse categorical cross-entropy (SCCE) are two additional variations of the Cross-Entropy. While both are used for multiclass classification, the latter is optimised for datasets with integer encoded labels, offering computational efficiency benefits. One challenge of using CE is that it can be sensitive to class imbalance, where one class has significantly more samples than the other. This can lead to biased models.
@@ VI-C2 Cross-Entropy (CE)
where $i$ iterates over all possible classes, $p_{i}$ is the learned probability of class $i$, and $q_{i}$ is the true posterior distribution of class $i$.In the context of AE end-to-end learning, the learned probability distribution is defined $p_{\theta_{r}}(a|y)$ as shown in Fig. 4 (b). The encoder with the learning parameter $\theta_{t}$ is trained to output the symbol distribution $f_{\theta_{t}}(a)$ based on the input data $a$. The optimisation of the AE model is then promised by minimising CCE [36] as
@@
TABLE III: Loss Functions Comparison
Loss
Function
Advantages Disadvantages Use Cases
MSE
• Intuitive: measures the average
squared difference between predicted and true values.
• Widely used in regression tasks.
• Sensitive to outliers: outliers can have
a significant impact on the loss.
• Less suitable for classification tasks:
may not be as informative for
classification problems.
Regression tasks
BCE
• Interpretable: Directly measures dissimilarity
between predicted and true distributions.
• Differentiable: suitable for gradient-based optimisation.
• Efficient: can be computed efficiently.
• Widely used: standard loss function in many
ML frameworks.
Sensitive to class imbalance
and assumes independence of classes
Binary
classification tasks
CCE Same as BCE Same as BCE
Multi-class
classification tasks
SCCE
More efficient for large datasets with
integer-encoded labels
Same as binary CE
Multi-class
classification tasks
with integer-encoded
labels
VI-D Physical constraints
@@
The constraints can be realised in several ways. One way is to incorporate a regularisation layer into the trainable network, within the AE encoder part. This layer can effectively maintain the level of encoded signals by amplitude $|w_{i}|\leq 1\ \forall w_{i}$, average power $\mathop{\mathbb{E}}[|w_{i}|^{2}]\leq 1\ \forall w_{i}$, or energy $||w||_{2}^{2}\leq n$, where $n$ is channel use [38, 39]. Some constraints such as reducing PAPR and ACL can be overcome by introducing an additional loss function that penalises the AE encoder to promote a restriction on the transmit signal such as [40].
@@
where $\mathcal{L}_{AE}$ is the AE loss function like MSE or CE, and $\mathcal{L}_{add}$ is the added penalties to realize the constraints. $\alpha_{AE}$ and $\alpha_{add}$ are the weights in penalties that help determine how much importance is given to minimising the loss of the learning process. This penalty is applied exclusively to the encoder weights[24, 25, 41].
VI-E AE Architecture Variant
@@
VI-E1 Classical AEs
They typically use fully connected layers (also known as dense layers) [19]. This is the simplest form of an AE and is designed for general-purpose dimensionality reduction or feature learning tasks. In this architecture, both the encoder and the decoder are composed of fully connected layers, where each neurone in a layer is connected to every neurone in the subsequent layer. This means that the AE maps the input directly to a lower-dimensional latent space and back to the original space.
VI-E2 Convolutional AEs
They are based on convolutional neural network (CNN), the convolutional AEs utilise convolutional layers instead of fully connected layers. This architecture is more suitable for handling large, high-dimensional data by efficiently learning spatial patterns. Several CNN architectures have been proposed, such as AlexNet, ZefNet, ResNet, etc., more details are available under [19].
VI-E3 Recurrent AEs
@@ VI-E4 Variational AEs
They are a generative model that learns the input data distribution and can generate new data samples [42]. Unlike classical AEs that learn a deterministic mapping from input to latent space, variational AEs (VAEs) learn a probabilistic latent space. This makes them more robust for generating realistic and diverse outputs. The encoder in VAEs outputs two vectors: the mean $\mu$ and the standard deviation $\sigma$ of a latent variable distribution. Instead of encoding the input to a fixed latent vector, VAEs assume that the latent space follows a Gaussian distribution. In addition to the reconstruction loss function that measures how well the reconstructed data match the input, the Kullback-Leibler (KL) divergence loss is used to measure how close the learnt latent distribution is to the standard Gaussian distribution:
VI-E5 Sparse AEs

### v2

@@ V AE-based Communication systems
The main goal of a communication system, acting as an AE, is to find a robust latent space representation $w$ of the input data in the transmitter to counteract any impairments introduced by the channel. In particular, through an end-to-end training, the AE learns to effectively reconstruct the input data at the receiver, considering the presence of channel impairments. During the training process, the encoder component of the AE acquires resilient symbol sequence representations for all incoming data. Once training is complete, both the transmitter and the receiver can function independently, using their respective layer dimensions and weights. Next, we delve into the various aspects of employing AEs in communication systems, encompassing the two AE approaches, activation functions, loss functions, physical constraints, and AE architectural variants.
V-A AE Architecture Approaches
In the context of communication systems, there are two variant ways of dealing with the communication problem, either as a regression problem or a categorical problem [32]. A regression problem in the context of machine learning and communication systems typically involves predicting continuous or numerical data based on the input data. By setting the output layer of the AE to a regression layer (such as a linear layer), the AE can be trained to reconstruct a continuous target value, which can be used effectively in communication systems. In contrast, in classification problems, the predicting data is a categorical or discrete class. End-to-end AE learning for communication systems using a classification problem involves training the AE to recognize and classify different constellation diagrams such as QPSK, 16-QAM, and 64-QAM. This can be done by setting the output layer of the AE to a classification layer (such as the sigmoid activation layer) and compute the cross-entropy.
@@ V-A AE Architecture Approaches
Fig. 4 shows three different end-to-end learning architectures based on AE. Regression problems can be used in AEs for end-to-end learning of communication systems. By setting the output layer of the AE to a linear regression layer as shown in Fig. 4 (a), the AE can be trained to predict continuous values for new inputs[33, 34]. This results in complex output data consisting of both in-phase (I) and quadrature (Q) components. This data is then fed into a demapper for symbol detection and subsequently a hard decision circuit to recover the original bit sequence. The AE encoder processes complex baseband symbols ($x\in\mathbb{C}$) from modulation formats such as quadrature phase-shift keying (QPSK), or M-quadrature amplitude modulation ($M$-QAM). The encoder, parametrized by $\theta_{t}$, generates resilient latent representations ($w\in\mathbb{C}$) by mapping $f_{\theta_{t}}:x\mapsto w$. These representations are then passed through a channel model, where they undergo impairments. The decoder, parameterized by $\theta_{r}$, aims to reconstruct the original symbol from the channel-corrupted output ($y\in\mathbb{C}$) by mapping $f_{\theta_{r}}:y\mapsto\hat{x}$, where $\hat{x}\in\mathbb{C}$. As it tries to predict the data of a numerical symbol, the Mean Squared Error (MSE) is a suitable loss function; see Section V-C. Since NNs cannot directly process complex data, complex input data $N$ must be converted to $2N$ real-valued representation for both the encoder and decoder by implementing a $\mathbb{C}2\mathbb{R}$ layer. The resulting real-valued output is then converted back to complex data using a $\mathbb{R}2\mathbb{C}$ layer.
Symbol-wise AEs and Bit-wise AEs are two different approaches to designing a communication system. Their primary difference lies in how they handle the mapping of bits to symbols and bit-labeling to effectively determine the learned constellation. This may directly impact system performance characteristics such as bit error rate (BER), symbol error rate (SER), and computational complexity. In both symbol-wise AE and bit-wise AE, the input to the encoder AE can be pre-processed to one hot encoded.
@@
As shown in Fig. 4 (b), a symbol-wise approach based on minimizing categorical cross-entropy (CCE) can be fed with a message $a\in\{0,1,...,M-1\}$, $M$ denotes the number of different symbols as $M=2^{m}$ where $m$ is the number of bits per symbol. Thus, the encoder then processes the mapping to complex data as $f_{\theta_{t}}:\{0,1,...,M-1\}\mapsto w$, i.e. $f_{\theta_{t}}:a\mapsto w$. Another key distinction lies in the output layer, the Softmax activation layer, which is used to obtain a probability vector of length $M$ which corresponds to the probabilities of possible $M$ messages, $p_{\theta_{r}}(a|y)$ [14]. This represents the estimated posterior distribution of the transmitted signal given the received signal. The constellation obtained through symbol-wise training is not optimized for bit-labeled. Bit-labeled indicates that each constellation point must be uniquely mapped to a specific bit vector. To enable joint constellation optimization and bit labeling, a bit-wise AE can be used, as shown in Fig. 4 (c), the bit vector $b$ can be fed to the encoder to perform the mapping $f_{\theta_{t}}:b\mapsto w$, where the mapping of bits into symbols is integrated within the encoder. The AE output is used directly without applying an activation function. These output values are called logits ($l$) which is a raw data and represents un-normalized probabilities of the data belonging to a specific class. Logits are normally suitable for classification tasks. Therefore, the loss function of binary cross-entropy (BCC) is suitable for such an approach. The Sigmoid function outputs a probability for each bit independently, which is applied to the decoder logits $l$ . This allows for the direct computation of the estimated posterior probabilities $p_{\theta_{r}}(b|y)$ for each bit, facilitating the use of log-likelihood ratios (LLRs) in the analysis [14]. The BCC minimization function can then be used to optimize AE, more details are given in Section V-C.
Although symbol-wise AEs suit low-complexity applications, bit-wise AEs are preferred for high-performance systems where joint optimization is critical, albeit at increased computational cost. The decision should weigh trade-offs between the complexity of training, the performance of the system, and the feasibility of the hardware.
V-B Activation Functions
@@
In the context of using AE for end-to-end learning in communication systems. The ReLU is widely used in hidden layers of encoder and decoder parts of the AE model. The ELU activation function is a non-saturating function that accelerates training while mitigating the risk of gradient explosion or vanishing gradient problems. The existing literature commonly employs ReLU and/or ELU activation functions in the hidden layer of AE-based communication systems. For example, [35] initially employed ReLU in their AE model but later switched to ELU for better results. Therefore, it is highly recommended to investigate both in order to determine the optimal choice. Furthermore, further research is needed to determine the combination use of both, for example, one for the encoder and the other for the decoder. However, the LeakyReLU activation function allows for a small, non-zero gradient when the input is negative, which helps mitigate the ”dying ReLU” problem often encountered with standard ReLU activations. This ensures a small gradient when the unit is inactive, which helps to keep the learning process alive [36]. Tanh activation has zero-centered output, which can make optimization easier. It is also useful in hidden layers to model complex relationships. In [37], Tanh activation is used to extract the feature. In [38], it is used in all the 1D convolutional layers. The scaled exponential linear unit (SELU) can induce self-normalization, which means that the activation automatically converges to zero mean and unit variance. This helps maintain the stability of the model during training, which is crucial for AEs that require consistent encoding and decoding [39]. It is suited for the hidden layer of AEs to leverage its self-normalization, which can lead to faster convergence. More research is required to investigate the potential benefits of alternative activation functions, such as Swish, for better gradient flow in the hidden layers. Finally, Softmax and Sigmoid activation layers are typically used in the output layer for classification tasks. The first can be used for symbol detection, while the second can be used for binary detection.
V-C Loss Functions
V-C1 Mean Squared Error (MSE)
@@
It serves as a loss function during the training phase, which is determined by computing the mean squared deviations between the predicted values and the actual ground truth values [40]. The general expression of MSE is
@@
Gradient-based learning [41, 42] is used to find optimal parameters $\theta=\{\theta_{t},\theta_{r}\}$ that minimize the loss function $\mathcal{L}_{MSE}(\theta)$ as
@@
where $\beta$ represents the learning rate, and the stochastic gradient Descent, such as Adam optimizer, enhances convergence.
V-C2 Cross-Entropy (CE)
CE is a widely used metric for training classification models as it is differentiable, which is crucial for gradient-based optimization algorithms[43, 44]. It is a quantitative metric in information theory that is used to measure the divergence between two probability distributions. This means that it measures the dissimilarity between two probability distributions. In the context of neural networks, one distribution is the output of the predicted probability distribution by the network, and the other is the true (target) distribution. Categorical cross-entropy (CCE) and sparse categorical cross-entropy (SCCE) are two additional variations of the Cross-Entropy. While both are used for multiclass classification, the latter is optimized for datasets with integer encoded labels, offering computational efficiency benefits. One challenge of using CE is that it can be sensitive to class imbalance, where one class has significantly more samples than the other. This can lead to biased models.
@@ V-C2 Cross-Entropy (CE)
where $i$ iterates over all possible classes, $p_{i}$ is the learned probability of class $i$, and $q_{i}$ is the true posterior distribution of class $i$.In the context of AE end-to-end learning, the learned probability distribution is defined $p_{\theta_{r}}(a|y)$ as shown in Fig. 4 (b). The encoder with the learning parameter $\theta_{t}$ is trained to output the symbol distribution $f_{\theta_{t}}(a)$ based on the input data $a$. The optimization of the AE model is then promised by minimizing CCE [45] as
@@
TABLE II: Loss Functions Comparison
Loss Func. Advantages Disadvantages Use Cases
MSE averages squared error between predictions and actual values Sensitive to outliers and Less suitable for classification tasks Regression tasks
BCE quantifies prediction errors against actual distributions, differentiable Sensitive to class imbalance and assumes independence of classes classification tasks
CCE Same as BCE Same as BCE Multi-class classification
SCCE optimized for large datasets with numeric labels Same as binary CE same as CCE but with integer labels
V-D Physical constraints
@@
The constraints can be realized in several ways. One way is to incorporate a regularization layer into the trainable network, within the AE encoder part. This layer can effectively maintain the level of encoded signals by amplitude $|w_{i}|\leq 1\ \forall w_{i}$, average power $\mathop{\mathbb{E}}[|w_{i}|^{2}]\leq 1\ \forall w_{i}$, or energy $w||_{2}^{2}\leq n$, where $n$ is channel use [13, 47]. Some constraints such as reducing PAPR and ACL can be overcome by introducing an additional loss function that penalizes the AE encoder to promote a restriction on the transmit signal such as [48].
@@
where $\mathcal{L}_{AE}$ is the AE loss function like MSE or CE, and $\mathcal{L}_{add}$ is the added penalties to realize the constraints. $\alpha_{AE}$ and $\alpha_{add}$ are the weights in penalties that help determine how much importance is given to minimizing the loss of the learning process. This penalty is applied exclusively to the encoder weights[35, 49].
V-E AE Architecture Variant
@@
V-E1 Classical AEs
They typically use fully connected layers (also known as dense layers) [50]. This is the simplest form of an AE and is designed for general-purpose dimensionality reduction or feature learning tasks. In this architecture, both the encoder and the decoder are composed of fully connected layers, where each neuron in a layer is connected to every neuron in the subsequent layer. This means that the AE maps the input directly to a lower-dimensional latent space and back to the original space.
V-E2 Convolutional AEs
They are based on convolutional neural network (CNN), the convolutional AEs utilize convolutional layers instead of fully connected layers. This architecture is more suitable for handling large, high-dimensional data by efficiently learning spatial patterns. Several CNN architectures have been proposed, such as AlexNet, ZefNet, ResNet, etc., more details are available under [50].
V-E3 Recurrent AEs
@@ V-E4 Variational AEs
They are a generative model that learns the input data distribution and can generate new data samples [51]. Unlike classical AEs that learn a deterministic mapping from input to latent space, variational AEs (VAEs) learn a probabilistic latent space. This makes them more robust for generating realistic and diverse outputs. The encoder in VAEs outputs two vectors: the mean $\mu$ and the standard deviation $\sigma$ of a latent variable distribution. Instead of encoding the input to a fixed latent vector, VAEs assume that the latent space follows a Gaussian distribution. In addition to the reconstruction loss function that measures how well the reconstructed data match the input, the Kullback-Leibler (KL) divergence loss is used to measure how close the learned latent distribution is to the standard Gaussian distribution:
V-E5 Sparse AEs